OpenStudy (anonymous):

MEDAL If x varies inversely with y and x = 8 when y = 10, find y when x = 6. A. Y=3/2 B. Y=4.8 C. Y=7.5 D. Y=40/3 i suck at inverse variation and don't understand it so can you please explain your answer

OpenStudy (gorv):

varies inversely means x\[Y \alpha \frac{ 1 }{ x }\]

OpenStudy (gorv):

to remove proportionality sign let it is \[y=\frac{ k }{ x}\]

OpenStudy (gorv):

now we made a relation in x and y where k is constant now first x and y are given from there find k value

OpenStudy (gorv):

tell me what u got for k ??

OpenStudy (anonymous):

SO Y=K/6

OpenStudy (gorv):

after getting k we have x=6 so y will be \[y=\frac{ k }{ 6 }\]

OpenStudy (gorv):

first calculate k

OpenStudy (gorv):

given x= 8 and y=10 put these values and find k

OpenStudy (anonymous):

k so 80=k im sorry i rlly suck at this